OpenStudy (anonymous):

1. How many grams of CaCl2 would be required to produce a 3.5 M (molar) solution with a volume of 2.0 L? 2. What is the molarity of a 5.00 x 102 ml solution containing 2490 g of KI? 3. How many moles of LiF would be required to produce a 2.5 M solution with a volume of 10.5 L?

OpenStudy (anonymous):

1) Ca=40.08 Ci^2 =2(35.45)=70.9 so she did 40.08+70.9=110.98 =1mole of CaCl 3.5m=7/2.0L 7(110.98g)=776.86g

OpenStudy (anonymous):

2) K=39.10 l=126.90 39.10+126.90=166.00+166.00=75/5=15m 166.00(15)=2490

OpenStudy (anonymous):

3)Li=6.94+F=19.00 6.94+19.00=25.94g LiF=25.94g 2.5M/10.5L 2.5solution=26.25moles/10.5L

OpenStudy (anonymous):

might want to check these ;D